# Generative User-Experience Research for Developing Domain-specific Natural Language Processing Applications

### v4

@@ 1 Introduction
The proposed methodology combines data- and user-driven approaches to full-cycle UX-driven prototype development for NLP applications. Prototyping or developing the first minimal viable products (MVPs) most benefits from the methodology. The proposed methodology consists of the following phases. First, the methodology describes generative UX research steps to explore opportunities for ideas and application concepts. This phase relies on exploratory data analysis and contextual inquiries, i.e., user interviews and on-site observations, as a source of idea generation. Furthermore, the domain users validate a proposed concept and help define the most important features for prototype implementation. Second, the prototyping addresses the most important identified research challenges crucial for usability assessment. Lastly, the prototype evaluation measures user utility via contextual inquiries before and after interacting with the prototype. We provide a project timeline with UX- and NLP-specific stages and tasks as a template for developing domain NLP applications.
Our team comprises UX researchers, data scientists, software developers, and a project leader because such a team composition facilitates a deeper understanding of the capabilities and limitations of the domain NLP application [19]. Therefore, the paper may have specific interest for the individual practitioners with the abovementioned roles or similar teams. Moreover, practitioners of user-driven research, e.g., customer research departments and quality assurance officers, could benefit from the proposed methodology and reported results.
@@ 2 Related Work
User experience research has become a requirement for practical and effective software development [22, 37]. UX puts perspective onto users, thus defining functionality and look of new products or features and ensuring a need in new products [26]. UX research methods analyze how users interact with a system of interest and explore what users miss or find inconvenient. Quantitative (e.g., questionnaires) and qualitative (e.g., interviews) methods are used to collect and analyze user feedback [49]. Structured interviews with domain-specific probe questions help UX elicit knowledge about the domain essential for better understanding system requirements [32].
Domain-driven design (DDD) is a widely used methodology of software development that aims to reflect real world domains through the architecture and interaction within and with the system [43, 36, 53, 30, 31]. UX focuses on learning about domains and domain-specific tasks to identify actors, entities, and processes and formalize them as a preliminary step of software engineering.
@@ 3 Methodology
In this section, we propose a methodology of generative UX research for optimizing a data-driven domain-specific NLP application toward higher user satisfaction. Generative research encourages user-focused development by getting to know the users better and explores opportunities for an ML/NLP application by learning users’ stories, pain points, and needs [54, 4]. Although generative research has already been employed in design and software development, it lacks coverage of full-cycle development of ML/NLP applications. For example, [4] focused on how to learn about user needs and experiences. [29] explored how to create a product concept based on the learned information from users and employed the users to evaluate this concept. Similarly, design thinking proposes a user-centered process of iterative ideating and prototyping [9, 38, 18]. [13] employed users to evaluate a prototype before developing a product for the production-ready state. Moreover, these methodologies do not consider an aspect of data-driven research that should be considered when creating an NLP application. To enable the full-cycle prototype development, the proposed methodology combines these previous efforts on generative research, design thinking, and domain understanding through exploratory data analysis required in ML/NLP applications.
Figure 1: Project timeline for the proposed methodology of generative UX research for prototyping domain NLP applications used in our case study. The methodology consists of three stages: exploration (data exploration and generative research), prototyping (development of the prototype and evaluation (assessment of accuracy and user utility). The methodology combines UX and NLP tasks, encouraging the teams to collaborate closely, especially during the generative research phase. The timeline reflects the reported case study’s duration, but each phase’s duration may differ when applied to other projects.
The methodology for full-cycle prototype development through generative research combines UX+ML/NLP to (1) explore ideas for an NLP application based on the stories of domain users, (2) collect feedback on a concept of a proposed NLP solution to their identified needs or pain points, (3) conduct a user-driven evaluation of a prototype from perspectives of user experience with a new NLP solution and accuracy estimation. Figure 1 depicts the proposed methodology as a timeline we followed to develop and evaluate a domain-specific NLP application. The methodology consists of three steps:
@@
    Prototyping to develop a fully working prototype with back- and frontend.
@@ 4.1 Exploration
4.1.1 Domain understanding and exploratory data analysis
A goal of domain understanding and exploratory data analysis is to learn about the area for the prototype and investigate data caveats that can either support or limit various stages of development. Our domain exploration steps included qualitative analysis, i.e., examining text data and identifying major roles of the domain users, and quantitative analysis, i.e., statistical analysis of the text records and composition of domain vocabulary.
@@ 4.1.1 Domain understanding and exploratory data analysis
4.1.2 Idea generation through user interviews
The objective of generative research is to learn how domain users use the system daily and elicit ideas for an NLP-based solution to improve the effectiveness of their existing processes. We performed contextual inquiries, i.e., one-to-one interviews with representative domain users in their environment [6]. We organized a week-long workshop at a plant to ensure an immersed environment for the research team [51, 6].
User interviews
@@ 4.1.2 Idea generation through user interviews
We summarized the interviews as follows: (1) a timeline of a shift that generalizes three stories of the shift leaders, (2) an overview of answers to the last two questions from all interviewees. These forms of summarization followed two different goals. First, to find the patterns in the user stories and identify which challenges or problems the interviewees mention at each step that we can turn into ideas for an NLP application. Second, to get an overview of the diverse opinions that can enable a broader picture for generating ideas.
Figure 5: A generalization of a shift from a shift leader resulting from our interviews. The context inquiries identified three phases of a shift and the main tasks and events that may occur during the day. We provide a short description and outline challenges for the steps in all phases, which we can turn into opportunities for an NLP application.
@@
Figure 5 generalizes an average shift of a shift leader. We identify three phases of a shift, i.e., beginning, middle, and end of the shift that determine which tasks a shift leader needs to perform. For the steps in the phases, we indicated a motivation for a step and which mentioned aspects we could turn into opportunities. For example, at the beginning of a shift, a shift leader reads through the logs of the past 24 hours to collect an overview of what happened when they were not on duty. We learned that if not enough information was logged, they talked to their colleagues, and the information from the oral communication remained unrecorded. We noted that recording the additional information from communication was an opportunity to improve data completeness in the logging system. We identified multiple cases throughout a shift, which indicated that more data could be logged to assist in more informed daily operations. Additionally, we learned that shift leaders rarely use the system to research a solution for solving incidents. The most common way to find information of interest is to read through the recent records or use a string-match search to find specific keywords, e.g., product names or order IDs. This aspect would be an opportunity to facilitate knowledge sharing in the logging system.
@@
Concept proposal: idea for a prototype
Our goal for the NLP application was to improve the information exchange, i.e., logging and retrieving crucial information. We expected the effective and efficient information exchange to yield higher production efficiency and shorter plant downtime. From the interviews, we learned about the common problems of data completeness, varying text quality, and the short life span of the records. To align with this goal and address the identified problems, we proposed turning the logging system of daily operations into a source of shared domain knowledge (Figure 7). As of the current state, only the most recent logged information is important for the users. We assumed that an NLP-driven domain knowledge processing system would enable leveraging the collected knowledge over the long run. Ultimately, users should be encouraged to log more information and improve their domain expertise by learning from the logged experience.
Figure 7: The historical data of the plant operation contains a lot of know-how information on daily operations, but the use of it is quite limited due to the inefficient search. By implementing efficient semantic search and allowing users to find more complex records, the project aims at motivating the logging of more detailed descriptions of daily operations, thus improving the input data completeness and quality.
@@ 4.1.3 Moderated concept evaluation
A click-dummy prototype aims at following two users’ mental models: the logging system and commonly used search engines, e.g., Google. Figure 8 shows that the prototype contains the visual attributes of a search field, a list of the resulting records, and such additional features as highlighting the matching excerpts and a similarity score. Moreover, its overall visual design is familiar to the users by resembling the overall design of the logging system.
@@
We conducted a second round of interviews where we presented the click-dummy prototype and evaluated our concept via guided interviews and moderated usability sessions. In the one-to-one interviews, we aimed to learn if providing efficient access to historical data would allow perform daily tasks in a more informed and effective manner:
@@
All interviewees agreed that efficient search would enable them to access the previously logged knowledge, e.g., solutions to problems. Moreover, they highlighted that a search application would provide additional assistance on a night shift and could assist in planning product launches. The shared ideas for the queries helped us to create lists of queries for the later prototype evaluation (see Section 4.3). Figure 6, the last column, overviews the interviewees’ replies on how they would use a system of shared domain knowledge. For example, they highlighted a need to access solutions to previously occurring problems. Moreover, they expected to retrieve information on a level of deep understanding. For example, for a query “quality state of product_A on Monday,” they would expect to see information about the batch level on that day. Additionally, they would explore information in an aggregating way, e.g., assemble problems that occurred to a specific product.
@@ 4.2 Prototyping
4.2.1 Pipeline
The search engine consists of two processes: (1) indexing of text records from a SQL database, (2) retrieving relevant records to a given query [^2] We use the terms “documents” to refer to the records of the logging system, i.e., short documents.. In the first prototype, we address a problem of the domain-specific language by using context expansion (i.e., adding definitions of some domain terms to the documents) and encoding documents with fastText (i.e., vector word representation capable of handling out-of-vocabulary domain words).
@@ 4.2.1 Pipeline
Indexing starts with accessing an SQL database and reading values from string columns (Figure 9). We concatenate all text fields of each record into one text document. Preprocessing performs context expansion to add semantic descriptors to some domain terminology (Section 4.2.2). We transform documents into multiple vector spaces: (1) vectorization into boolean and term frequency-inverse document frequency (TF-IDF) representations that require full-text collection, (2) per-document vectorization into semantic representation, which employs word embedding service to represent words into fastText word vectors and computes TF-IDF weighted-mean word vectors to represent documents. Finally, we save boolean and semantic document representations into a search index. Document retrieval uses the same word embedding service and represents a query as a mean of word vectors (Figure 10). First, candidate retrieval retrieves documents that contain terms to be mandatory present in the documents (depending on if a query contains terms in quotes). Second, it filters the documents by semantic similarity (Section 4.2.4). Lastly, we rank the top M most semantically similar documents by combined semantic similarity on the document and term levels.
4.2.2 Preprocessing
@@
4.2.3 Representation of the domain-specific language
@@ 4.2.4 Retrieval and ranking
4.2.5 User interface
The user interface of our prototype resembles a click-dummy prototype proposed for the concept evaluation (see Section 4.1.3). Figure 12 depicts an interface of the prototype that contains the text of the record, enables sorting by relevance and timestamp, and contains functionality to collect user feedback for the prototype evaluation. Additionally, the prototype provides a detailed view by clicking on a record of interest. For the first version of the prototype, we excluded the functionality of highlighting and visualization of the relevance score to prevent judgment bias during the user-based evaluation of the prototype accuracy (see Section 4.3.1).
@@ 4.3 Evaluation
Our user-driven prototype evaluation aims to assess prototype accuracy and how user utility and effectiveness change after the open interaction with the prototype. We recruited a focus group of seventeen domain experts and advanced users of the logging system from one plant for the prototype evaluation with a diverse set of roles to ensure opinion variance in the user studies (Table 1). Some study participants were previously recruited for the exploratory generative UX research (see Section 4.1.2). The users were asked to perform two types of interaction: (1) perform a set of tasks to assess the relevance of search results, and (2) freely use the system as they see fit during the execution of their work tasks. Fourteen participants who could perform tasks for the relevance assessment were invited to fill in a questionnaire before and after the execution of the task. All participants were encouraged to use the system on demand and were invited to the interviews.
@@
Role Invited to the prototype evaluation Relevance assessment & Questionnaires Interviews
@@
4.3.1 Relevance assessment
The relevance assessment focused on the precision-oriented IR evaluation and estimated the relevance of the $N$ top-ranked results. Although such “shallow” evaluation could be incomplete, it is consistent if users are interested only in top-ranked documents in the domain application [39]. Therefore, the relevance assessment enabled the estimation of the prototype accuracy in daily use.
Experiment setup: Collecting a test dataset
Document relevance level # queries $\kappa$
term level 44 .856
phrase level 16 .673
Methodology
Baselines
After implementing the prototype, we provided two baselines to evaluate how user utility changed: a keyword search (i.e., exact term match) and BM25. The keyword search is a default SQL-based search from the logging system with which the domain users are familiar. BM25 represents a standard baseline used in IR.
We implemented a more flexible version of the keyword search for the experiments. We retrieved a document if at least one term from a query was found. The documents were ranked by the decreasing number of overlapping terms and then by a timestamp of the records.
For BM25, we normalized German text records and accommodated local domain setup with lemmatizing and removing stopwords. We retrieved documents if cosine similarity was higher than 0.15 and ranked them using it and a document’s timestamp.
We implemented an optional context expansion in both baselines to evaluate its impact on handling the domain-specific language (see Section 4.2.2).
Metrics
We evaluated the accuracy of the prototype at top $N$ records with precision-oriented metrics, i.e., precision $P$@N, and metrics for ranking quality, i.e., mean average precision $mAP$@N [39], mean reciprocal rank $MRR$ [47] and normalized discounted cumulative gain $nDCG$@N [34]. We calculated metrics for the top 5 and 20 records in our experiments. The combined relevance scores in the test dataset were used to compute $nDCG$@N.
Results
Using context expansion improved precision $P$@5 and $P$@20 of both baselines. Figure 14 depicts the influence of context expansion. The ranking values measured by $MRR$, $mAP$@5, and $mAP$@20 that ignore the annotated relevance score of the documents have slightly dropped. On the contrary, $nDCG$@5 and $nDCG$@20 that consider the relevance score have increased. The results show that context expansion assisted in retrieving more candidate documents and ranking relevant documents higher.
keyword search - 3300 .153 .068 .125 .076 .063 .112 .096
+ 5944 .132 .180 .077 .152 .101 .066 .100
BM25 - 736 .582 .294 .548 .322 .178 .479 .291
+ 657 .572 .302 .523 .337 .182 .459 .291
prototype: semantic search + 90 .888 .749 .880 .725 .668 .805 .772
4.3.2 UX evaluation
UX evaluation estimated how user utility and effectiveness change before and after using the prototype, e.g., if users found the information more efficiently and effectively. Additionally, through interviews, we collected free-form feedback on using the prototype and ideas for improvements from a user perspective.
Questionnaires

### v5

@@ 1 Introduction
The proposed methodology combines data- and user-driven approaches to full-cycle UX-driven prototype development for NLP applications. Prototyping or developing the first minimal viable products (MVPs) most benefits from the methodology. The proposed methodology consists of the following phases. First, the methodology describes generative UX research steps to explore opportunities for ideas and application concepts. This phase relies on exploratory data analysis and contextual inquiries, i.e., user interviews and on-site observations, as a source of idea generation. Furthermore, the domain users validate a proposed concept and help define the most important features for prototype implementation. Second, the prototyping addresses the most important identified research challenges crucial for usability assessment. Lastly, the prototype evaluation measures user utility via contextual inquiries before and after interacting with the prototype.
Our team comprises UX researchers, data scientists, software developers, and a project leader because such a team composition facilitates a deeper understanding of the capabilities and limitations of the domain NLP application [19]. Therefore, the paper may have specific interest for the individual practitioners with the above-mentioned roles or similar teams. Moreover, practitioners of user-driven research, e.g., customer research departments and quality assurance officers, could benefit from the proposed methodology and reported results.
@@ 2 Related Work
Domain-driven design (DDD) is a software development methodology that understands, conceptualizes, and models complex domains into applications closely aligned with business needs [22]. The core of DDD lies in close interaction between developers, stakeholders, and users to study and implement the business logic of the underlying processes. According to DDD, the building blocks of software should reflect the physical world as closely as possible in terms of both functionality and naming (e.g., [44, 37, 54, 31, 32]). The methodology emphasizes understanding a domain by understanding and utilizing its ubiquitous language by all development participants in all planning, development, and refinement stages. Moreover, DDD enables dynamic adjustments of a domain model to changes in real life through an iterative process of collecting feedback from domain users.
User experience research has become a requirement for practical and effective software development in addition to the domain understanding [23, 38]. UX puts perspective onto users, thus defining functionality and look of new products or features and ensuring a need in new products [27]. UX research methods analyze how users interact with a system of interest and explore what users miss or find inconvenient. Quantitative (e.g., questionnaires) and qualitative (e.g., interviews) methods are used to collect and analyze user feedback [50]. Structured interviews with domain-specific probe questions help UX elicit knowledge about the domain essential for better understanding system requirements [33].
@@ 3 Methodology
In this section, we propose a methodology of generative UX research for optimizing a data-driven domain-specific NLP application toward higher user satisfaction. Generative research encourages user-focused development by getting to know the users better and explores opportunities for an ML/NLP application by learning users’ stories, pain points, and needs [55, 4]. The methodology is grounded on a domain-driven design (DDD) [22] of deep domain understanding through close collaboration between developers and domain experts. The iterative knowledge exchange about domain know-how, problem-solving requirements, and implementation constraints leads to effective domain modeling, i.e., a conceptual representation of the domain at consideration. An effective and knowledge-rich domain model reflects internal and external business processes most fully. Moreover, a model should be clear to both domain experts to provide feedback and developers to act as a foundation for implementation. The generative research puts domain experts even more into perspective by switching from business processes to domain experts’ workflows and their daily routines.
Although generative research has already been employed in design and software development, it lacks coverage of full-cycle development of ML/NLP applications. For example, [4] focused on how to learn about user needs and experiences. [30] explored how to create a product concept based on the learned information from users and employed the users to evaluate this concept. Similarly, design thinking proposes a user-centered process of iterative ideating and prototyping [9, 39, 18]. [13] employed users to evaluate a prototype before developing a product for the production-ready state. Moreover, these methodologies do not consider an aspect of data-driven research that should be considered when creating an NLP application. The proposed methodology combines these previous efforts on generative research, design thinking, and domain understanding through exploratory data analysis required in ML/NLP applications.
Figure 1: A timeline for the proposed methodology of generative UX research for prototyping domain NLP applications used in our case study. The methodology consists of three stages: exploration (data exploration and generative research), prototyping (development of the prototype and evaluation (assessment of accuracy and user utility). The phases show the approximate share of the workload of the UX and NLP teams. For the majority of the phases, both teams need to be involved. Especially during the generative research phase, an equal involvement of the two teams enables deep domain understanding.
The methodology for full-cycle prototype development through generative research combines UX+ML/NLP to (1) explore ideas for an NLP application based on the stories of domain users, (2) collect feedback on a concept of a proposed NLP solution to their identified needs or pain points, (3) conduct a user-driven evaluation of a prototype from perspectives of user experience with a new NLP solution and accuracy estimation. Figure 1 depicts our proposed methodology to develop and evaluate a domain-specific NLP application. The methodology consists of three steps:
@@
    Prototyping to develop a fully working prototype with back- and front-end.
@@ 4.1 Exploration
4.1.1 Exploratory data analysis
A goal of domain understanding and exploratory data analysis is to learn about the area for the prototype and investigate data caveats that can either support or limit various stages of development. Our domain exploration steps included qualitative analysis, i.e., examining text data and workflows of domain experts, and quantitative analysis, i.e., statistical analysis of the text records and composition of domain vocabulary.
@@ 4.1.1 Exploratory data analysis
4.1.2 Workshop: Generative research
The objective of generative research is to learn how domain users use the system daily and elicit ideas for an NLP-based solution to improve the effectiveness of their existing processes. We performed contextual inquiries, i.e., one-to-one interviews with representative domain users in their environment [6]. We organized a week-long workshop at a plant to ensure an immersed environment for the research team [52, 6]. The workshop’s goal was to better understand experts’ processes and workflows through observation and interaction with them [22]. The observed processes were to be organized into explanatory models to be later distilled into ideas for the prototype implementation.
Contextual inquiries: User interviews
@@ 4.1.2 Workshop: Generative research
We summarized the interviews as follows: (1) a timeline of a shift that generalizes three stories of the shift leaders, (2) an overview of answers to the last two questions from all interviewees. These forms of summarization followed two different goals. First, we need to find the patterns in the user stories and identify which challenges or problems the interviewees mention at each step that we can turn into ideas for an NLP application. Second, to get an overview of the diverse opinions that can enable a broader picture for generating ideas.
Figure 5: A generalization of a shift from a shift leader resulting from our interviews that acts as a domain model for our user study. The context inquiries identified three phases of a shift and the main tasks and events that may occur during the day. We provide a short description and outline the challenges of the steps in all phases, which we can turn into opportunities for an NLP application.
@@
Figure 5 generalizes an average shift of a shift leader. We identify three phases of a shift, i.e., beginning, middle, and end of the shift, which determine which tasks a shift leader needs to perform. For the steps in the phases, we indicated a motivation for a step and which mentioned aspects we could turn into opportunities. For example, at the beginning of a shift, a shift leader reads through the logs of the past 24 hours to collect an overview of what happened when they were not on duty. We learned that if not enough information was logged, they talked to their colleagues, and the information from the oral communication remained unrecorded. We noted that recording the additional information from communication was an opportunity to improve data completeness in the logging system. We identified multiple cases throughout a shift, which indicated that more data could be logged to assist in more informed daily operations. Additionally, we learned that shift leaders rarely use the system to research a solution for solving incidents. The most common way to find information of interest is to read through the recent records or use a string-match search to find specific keywords, e.g., product names or order IDs. This aspect would be an opportunity to facilitate knowledge sharing in the logging system.
@@
Concept development with a click-dummy prototype
Our goal for the NLP application was to improve the exchange of information, i.e., logging and retrieving crucial information. We expected the effective and efficient information exchange to yield higher production efficiency and shorter plant downtime. From the interviews, we learned about the common problems of data completeness, varying text quality, and the short life span of the records. To align with this goal and address the identified problems, we proposed turning the logging system of daily operations into a source of shared domain knowledge (Figure 7). As of the current state, only the most recent logged information is important for the users. We assumed that an NLP-driven domain knowledge processing system would enable leveraging the collected knowledge over the long run. Ultimately, users should be encouraged to log more information and improve their domain expertise by learning from the logged experience.
Figure 7: The historical data of the plant operation contains a lot of know-how information on daily operations, but the use of it is quite limited due to the inefficient search. By implementing efficient semantic search and allowing users to find more complex records, the project aims to motivate the logging of more detailed descriptions of daily operations, thus improving the completeness and quality of input data.
@@
A click-dummy prototype aims to follow two users’ mental models: the logging system and commonly used search engines, e.g., Google. Figure 8 shows that the prototype contains the visual attributes of a search field, a list of the resulting records, and such additional features as highlighting the matching excerpts and a similarity score. Moreover, its overall visual design is familiar to the users by resembling the overall design of the logging system.
@@
Moderated concept evaluation
We conducted a second round of interviews where we presented the click-dummy prototype and evaluated our concept via guided interviews and moderated usability sessions. In the one-to-one interviews, we aimed to learn if providing efficient access to historical data would allow performing daily tasks in a more informed and effective manner:
@@
All interviewees agreed that efficient search would enable them to access the previously logged knowledge, e.g., solutions to problems. Moreover, they highlighted that a search application would provide additional assistance on a night shift and could assist in planning product launches. The shared ideas for the queries helped us to create lists of queries for the later prototype evaluation (see Section 4.3). Figure 6, the last column, overviews the interviewees’ replies on how they would use a system of shared domain knowledge. For example, they highlighted a need to access solutions to previously occurring problems. Moreover, they are expected to retrieve information with a deep understanding. For example, for a query “quality state of product_A on Monday,” they would expect to see information about the batch level on that day. Additionally, they would explore information in an aggregating way, e.g., assemble problems that occurred to a specific product.
@@ 4.2 Prototyping
4.2.1 Planning
MVP requirements for the prototype development
From the exploration phase and workshop, we deduct the following main requirements for the prototype: The search engine of the prototype shall be designed in a typical IR process architecture: (1) indexing of text records from a SQL database, (2) retrieving relevant records to a given query [^2] We use the terms “documents” to refer to the records of the logging system, i.e., short documents.. The most challenging requirement for the MVP is to tackle the problem described in the domain-specific language and technical style of the text records. It should use context expansion (i.e., adding definitions and descriptions of technical codes to the documents) and encoding documents vector word representation capable of handling out-of-vocabulary domain words, e.g., fastText. Furthermore, the MVP is supposed to be seamlessly integrated into the main software application.
@@ 4.2.1 Planning
4.2.2 Development - Backend & Frontend
Indexing
Indexing starts with accessing an SQL database and reading values from string columns (Figure 9). We concatenate all text fields of each record into one text document. Preprocessing performs context expansion to add semantic descriptors to some domain terminology. We transform documents into multiple vector spaces: (1) vectorization into boolean and term frequency-inverse document frequency (TF-IDF) representations that require full-text collection, (2) per-document vectorization into semantic representation, which employs word embedding service to represent words into fastText word vectors and computes TF-IDF weighted-mean word vectors to represent documents. Finally, we save boolean and semantic document representations into a search index. Document retrieval uses the same word embedding service and represents a query as an average of word vectors. First, candidate retrieval retrieves documents that contain terms to be mandatory present in the documents (depending on if a query contains terms in quotes). Second, it filters the documents by semantic similarity (Section 4.2.2). Lastly, we rank the top M most semantically similar documents by combined semantic similarity on the document and term levels.
Preprocessing
@@ 4.2.2 Development - Backend & Frontend
Representation of the domain-specific language
@@
User interface
The user interface of our prototype resembles a click-dummy prototype proposed for the concept evaluation. Figure 12 depicts an interface of the prototype that contains the text of the record, enables sorting by relevance and timestamp, and contains functionality to collect user feedback for the prototype evaluation. Additionally, the prototype provides a detailed view by clicking on a record of interest. For the first version of the prototype, we excluded the functionality of highlighting and visualization of the relevance score to prevent judgment bias during the user-based evaluation of the prototype accuracy (see Section 4.3.2).
@@ 4.3 Evaluation
Our user-driven prototype evaluation aims to assess how users utilize the prototype in their work routine, i.e., how the effectiveness of their daily tasks changes after the unmoderated interaction with the prototype, and how accurately the prototype performs. We recruited a focus group of seventeen domain experts and advanced users of the logging system from one plant for the prototype evaluation with a diverse set of roles to ensure opinion variance in the user studies (Table 1). Some study participants were previously recruited for the exploratory generative UX research (see Section 4.1.2). The users were asked to perform two types of interaction: (1) freely use the system as they see fit during their work tasks (user-driven prototype evaluation), (2) perform a set of tasks to assess the relevance of search results (accuracy evaluation).
@@
Role Invited to the prototype evaluation Performed relevance assessment & Questionnaires Participated in interviews
@@
4.3.1 User-driven prototype evaluation
UX evaluation estimated how user utility and effectiveness change before and after using the prototype, e.g. if users found the information more efficiently and effectively. The users were asked to complete a questionnaire and were interviewed before the unmoderated interaction with the prototype (i.e., control point). After several weeks of free use of the prototype, we repeated the questionnaires and interviews to reflect on how their pattern of acquiring information has changed. Additionally, through interviews, we collected free-form feedback on using the prototype and ideas for improvements from a user perspective. Below, we report quantitative analysis on the questionnaires and qualitative analysis of the interviews.
Questionnaires: Control point and reflection